OpenStudy (anonymous):

3x + 9y = 18 slope intercept ? :')

OpenStudy (jdoe0001):

if you solve " 3x + 9y = 18" for "y" what does it give you?

OpenStudy (anonymous):

I know the Y-intercept is 2 if that's what you're asking ?

OpenStudy (jdoe0001):

well.. nope... just to solve it for "y", a plain linear simplification :)

OpenStudy (anonymous):

Slope intercept form is y = mx + b where slope is m and the y-intercept is b Try converting the equation to where is fits this form! :)

OpenStudy (anonymous):

so.. y = -1/3x+2 .. ? idk what im doing sorry

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Ahh thanks you two ! c: <33

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

:D It's simple really. Good luck with future questions and you're welcome!